OpenStudy (anonymous):

Movies often use models to create special effects. In the movie Son of Godzilla, Godzilla appears to be 67 feet tall and was played by a man in a 7.5 ft rubber suit. Godzilla’s son appeared to be 45 feet tall in the movie; how tall was the rubber suit for the actor playing the son? (Round to the nearest hundredth)

OpenStudy (studygurl14):

Okay, so you will use proportions and ratios to find the answer. Godzilla: 67 ft Man playing him: 7.5 ft Godzilla's son: 45 Actor: ? \[\frac{ 67 }{ 7.5 } = \frac{ 45 }{ A }\] Cross multiply \[67A = 7..5(45)\] Now just solve for A. That will be your answer. Hope this helps!

OpenStudy (anonymous):

thanks

OpenStudy (studygurl14):

you're welcome